James works two part time jobs. one pays $7/hr and the other pays $10/hr. Last week he worked a total of 32 hours and made a total of $278. How many hours did he work at each job?
Answer by checkley77(12569)

10X+7(32-X)=278
10X+224-7X=278
3X=278-224
3X=54
X=54/3
X=18 HOURS WORKED AT THE $10 PER HOUR JOB.
32-18=14 HOURS WORKED AT THE $7 PER HOUR JOB.
PROOF:
10*18+7*14=278
180+98=278
278=278